# Inverse problems with experiment-guided AlphaFold

### v1

@@ A.3 Calculating $F_{\mathrm{c}}$ from a protein structure
A.4 Filtering & relaxation after density guidance
@@
A.5 Ensemble pruning with matching pursuit
@@
Before selecting the ensemble, we optimize a scalar B-factor $B$ to maximize the log-likelihood in equation (1). This step adjusts the bandwidth of the ensemble’s theoretical electron density $F_{\mathrm{c}}$ to best fit the observed density $F_{\mathrm{o}}$. While optimizing $B$ during the diffusion process can introduce numerical instabilities, we avoid the issue here because $B$ is optimized after structure generation, filtering, and relaxation is complete. Hence, we do not encounter similar instabilities here. We optimized $B$ using Adam (Diederik, 2014) optimizer with a step size of $1.0$ over $100$ iterations. The optimized B-factor $B^{\ast}$ is used uniformly across all atoms in the remaining structures. Following this, we apply the matching pursuit algorithm to select the best-fit ensemble.
Across all experiments, we set the maximum ensemble size to $n_{\text{max}}=5$. For NMR guided ensembles, we do not employ ensemble selection and our results are evaluated on the full ensemble.
A.6 Electron density pre-processing
@@
A.7 Ensemble bi-modal distribution evaluation
Figure A1: Distribution of normalized distances to conformations A ($-1$) and B (+1)+1)+ 1 ) in generated ensembles for $12$ protein structures. Compared are ensembles generated by AlphaFold3 and our electron density i.i.d. and non-i.i.d. guidance. While unguided AlphaFold3 typically fails to capture the multi-modal nature of these structures, non-i.i.d. consistently produces multi-modal ensembles.
@@ A.7 Ensemble bi-modal distribution evaluation
In this section, we describe a quantitative evaluation demonstrating that non-i.i.d. density guidance better captures the bimodality inherent to altloc regions compared to other methods. To this end, we use samples from three sets of experiments, density-guided (i.i.d.), density-guided (non-i.i.d.), and AlphaFold3 (unguided). In each case, we filter, relax (Appendix A.4), and refine the generated ensemble using matching pursuit-based selection (Appendix A.7). To quantify bimodality, we compute the normalized distance of each sample in the ensemble $\bm{\mathcal{X}}$ relative to the known altloc configurations (A or B). Consider an individual sample $\mathbf{X}\in\mathbb{R}^{m\times 3}$ from the ensemble $\bm{\mathcal{X}}$ and the reference altloc A and B structures $\mathbf{X}_{\text{a}},\mathbf{X}_{\text{b}}\in\mathbb{R}^{m\times 3}$, respectively. Following (Rosenberg et al., 2024b), we define the signed normalized distance between $\mathbf{X}$ and the reference conformations as
@@
This metric provides an interpretable method for quantifying ensemble distributions. Specifically, samples closer to reference conformation A have the normalized distance approaching $-1$, while those closer to reference conformation B have the normalized distance approaching $+1$. Consequently, negative values indicate proximity to mode A, while positive values indicate proximity to mode B.
The resulting normalized distance distributions are visualized in Figure A1. We notice that non-i.i.d. guided sampling consistently achieves bimodal and multimodal behavior with proximity to both reference conformations (positive and negative modes in the plot). In comparison, i.i.d. guided sampling achieves a moderate degree of bimodality but is less effective than non-i.i.d. sampling in terms of accurately capturing the full extent of conformational heterogeneity present in the electron density. In comparison, AlphaFold3 (unguided) often fails to correctly recover bimodal modal behavior. This demonstrates that non-i.i.d. guidance significantly outperforms i.i.d. guidance in modeling bimodal distributions in electron density maps. In addition, both guided approaches perform significantly better than AlphaFold3.
A.8 Diffusion guidance landscape
@@
Figure A2: Quantiative assessment of generated structure validity (top plot) and agreement with the experimental observation (bottom plot) during the progress of diffusion iterations using an electron density likelihood for guidance. Structural validity is quantified as the number of violated bond lengths, while the agreement to experiment is measured using the cosine similarity between $F_{\mathrm{o}}$ and $F_{\mathrm{c}}$.
@@
A.9 Estimating NOE restraints from an ensemble
@@
A.10 NMR data collection and pre-processing
@@
A.11 Computational estimation of N-H bond order
@@
A.12 Evaluation metrics used in NMR structure determination experiments
In what follows, we describe how we compute quantitative metrics reported in Table A4 and Figure 7. Given a structural ensemble $\bm{\mathcal{X}}$ and a restraint list $D=\{(\underline{d}_{ij},\overline{d}_{ij}):(i,j)\in\mathcal{P}\}$, we measure the quality of $\bm{\mathcal{X}}$ in terms of its adherence to $D$ and as a secondary measure, we evaluate the recovered conformational flexibility with respect to the NMR ensemble $\bm{\mathcal{X}}_{\text{NMR}}$. To evaluate restraint violation across the ensemble, we first compute the ensemble distance matrix $D({\bm{\mathcal{X}}})$, whose entries are given by $d_{ij}(\bm{\mathcal{X}})=\frac{1}{n}\sum_{k=1}^{n}d_{ij}(\mathbf{X}^{k})$.

### v2

@@ A.6 Calculating $F_{\mathrm{c}}$ from a protein structure
A.7 Filtering & relaxation after density guidance
@@
A.8 Ensemble pruning with matching pursuit
@@
Before selecting the ensemble, we optimize a scalar B-factor $B$ to maximize the log-likelihood in equation (1). This step adjusts the bandwidth of the ensemble’s theoretical electron density $F_{\mathrm{c}}$ to best fit the observed density $F_{\mathrm{o}}$. While optimizing $B$ during the diffusion process can introduce numerical instabilities, we avoid the issue here because $B$ is optimized after structure generation, filtering, and relaxation is complete. Hence, we do not encounter similar instabilities here. We optimized $B$ using Adam (Diederik, 2015) optimizer with a step size of $1.0$ over $100$ iterations. The optimized B-factor $B^{\ast}$ is used uniformly across all atoms in the remaining structures. Following this, we apply the matching pursuit algorithm to select the best-fit ensemble.
Across all experiments, we set the maximum ensemble size to $n_{\text{max}}=5$. We heuristically found that the density is well explained by at most $5$ samples and adding more samples to the ensemble would overfit the noise in the density map without yielding a considerable increase in cosine similarity. In Figure A6, we plot the normalized cosine similarity against the number of samples in an ensemble of size $16$. The cosine similarity plateaus at $5$ samples after which we either get a deteriorated fit to the density, or overfit to noise – both undesirable.
For NMR guided ensembles, we do not employ ensemble selection and our results are evaluated on the full ensemble.
A.9 Electron density pre-processing
@@
A.10 Ensemble bi-modal distribution evaluation
@@
In this section, we describe a quantitative evaluation demonstrating that non-i.i.d. density guidance better captures the bimodality inherent to altloc regions compared to other methods. To this end, we use samples from three sets of experiments, density-guided (i.i.d.), density-guided (non-i.i.d.), and AlphaFold3 (unguided). In each case, we filter, relax (Appendix A.7), and refine the generated ensemble using matching pursuit-based selection (Appendix A.10). To quantify bimodality, we compute the normalized distance of each sample in the ensemble $\bm{\mathcal{X}}$ relative to the known altloc configurations (A or B). Consider an individual sample $\mathbf{X}\in\mathbb{R}^{m\times 3}$ from the ensemble $\bm{\mathcal{X}}$ and the reference altloc A and B structures $\mathbf{X}_{\text{a}},\mathbf{X}_{\text{b}}\in\mathbb{R}^{m\times 3}$, respectively. Following Rosenberg et al. (2024b), we define the signed normalized distance between $\mathbf{X}$ and the reference conformations as
@@
This metric provides an interpretable method for quantifying ensemble distributions. Specifically, samples closer to reference conformation A have the normalized distance approaching $-1$, while those closer to reference conformation B have the normalized distance approaching $+1$. Consequently, negative values indicate proximity to mode A, while positive values indicate proximity to mode B. The resulting normalized distance distributions are visualized in Figure A2, A3, and A4. We observe that non-i.i.d. guided sampling with both Chroma and AlphaFold3 achieves bimodal and multimodal distributions behavior with proximity to both reference conformations (positive and negative modes in the plot). Notably, our electron density-guided approach with AlphaFold3 achieves bimodality more consistently than Chroma-based guidance. In contrast, i.i.d. guided sampling with AlphaFold3 achieves a moderate degree of bimodality but is less effective than non-i.i.d. sampling in terms of accurately capturing the full extent of conformational heterogeneity present in the electron density. Furthermore, unguided methods – including AlphaFold3, AlphaFlow, ESMFlow, and Chroma – frequently fail to recover the correct bimodal behavior.
This demonstrates that non-i.i.d. guidance significantly outperforms i.i.d. guidance in modeling bimodal distributions in electron density maps. In addition, both guided approaches perform significantly better than unguided models.
A.11 Diffusion guidance landscape
@@
A.12 Estimating NOE restraints from an ensemble
@@
A.13 NMR data collection and pre-processing
@@
A.14 Computational estimation of N-H bond order
@@
A.15 Evaluation metrics used in NMR structure determination experiments
In what follows, we describe how we compute quantitative metrics reported in Table A4 and Figure A7. Given a structural ensemble $\bm{\mathcal{X}}$ and a restraint list $D=\{(\underline{d}_{ij},\overline{d}_{ij}):(i,j)\in\mathcal{P}\}$, we measure the quality of $\bm{\mathcal{X}}$ in terms of its adherence to $D$ and as a secondary measure, we evaluate the recovered conformational flexibility with respect to the NMR ensemble $\bm{\mathcal{X}}_{\text{NMR}}$. To evaluate restraint violation across the ensemble, we first compute the ensemble distance matrix $D({\bm{\mathcal{X}}})$, whose entries are given by $d_{ij}(\bm{\mathcal{X}})=\frac{1}{n}\sum_{k=1}^{n}d_{ij}(\mathbf{X}^{k})$.